# Neural Post-Einsteinian Framework for Efficient Theory-Agnostic Tests of General Relativity with Gravitational Waves

### v1

@@ III.6 Training Procedure and Results
We implement the networks described in Secs. III.4 and III.5 with PyTorch [101]. For training, we choose a combination of the AdamW optimizer [102], with a weight decay of $10^{-4}$, and the ExponentialLR scheduler, with a learning rate decay of by a factor of $0.9$ every epoch. We start by training the VAE for $50$ epochs with a batch size of $64$ and an initial learning rate of $10^{-4}$. The coefficient of the KL loss is fixed to $\kappa=10^{-6}$ during the process. Then, we freeze the VAE parameters and train the secondary network for another 50 epochs using the same batch size and initial learning rate. Here, the value of $\kappa$ is empirically determined to give the best learning result for the shape function. The other settings, such as the learning rate and the batch size, are free to vary in some reasonable range, at most impacting the speed of convergence.
@@
Figure 4 shows the training and validation losses as functions of epochs for the VAE (panel (a)) and the secondary newtwork (panel (b)). Observe that both the VAE and the secondary network reach convergence within the maximum number of epochs explored. The validation loss of the secondary network is considerably higher than the training loss after the loss has converged, implying overfitting. However, as mentioned in Sec. III.5, the secondary network does not necessarily require high accuracy, since it only affects the EFT prior boundary. The validation loss after convergence implies a $\sim 6\%$ fractional error for the modeling of the scale function $T$ [^3] According to Eq. (31), this error is roughly the exponential of the square-root average loss minus one. Here, the average loss is read from the rightmost orange marker in Fig. 4 (b)., which is acceptable for our purposes.
@@
The representation of ppE theories in the latent space (given by the encoder mean) is shown in Fig. 5 (a). These ppE theories map to radial lines, which is not surprising given the npE prescription in Eqs. (14) and (15), and the fact that ppE phases associated with the same $b_{\mathrm{ppE}}$, do not vary in shape. What matters is that these ppE theories are well organized in the latent space. In particular, the mapping is injective, in the sense that there is a one-to-one relation between the ppE index $b_{\mathrm{ppE}}$ (or PN order) and the polar angle (up to $\pi$) of the latent line, so each theory maps to one line and different theories separate. Moreover, the mapping is ordered, in the sense that theories with deformations that enter at similar PN orders are also close to each other in angle in the latent space. We highlight that we arrive at this result in a completely unsupervised manner, i.e. ordering is not imposed explicitly.
@@
Figure 5 (a) also presents ppE theories that introduce modifications that enter at half-integer PN orders. These theories are not included in the modified gravity data set for training, but they are mapped to the latent space in a similar way to those in the training set. This means that the VAE understands well the structure of ppE theories in the latent space. These observations are further confirmed by Fig. 5 (b), where we present the shape functions generated from a sample of latent space angles, which is much denser than the distribution of angles of the ppE theories contained in the training set. These functions form a band when overlaid, suggesting that the ppE parameterization has been continuously extended by the npE latent space.
As discussed in Sec. III.4, the pseudo-PN expansion adopted by the decoder leads to a trivial extension of the ppE parameterization when $N_{p}=1$. Although we have chosen $N_{p}=2$, the two terms in Eq. (24) may lead to degeneracies after training, and reduce to the trivial case. Therefore, we must investigate the behavior of the learned pseudo-PN expansion across the latent space, which is illustrated in Fig. 6. In this figure, the blue (orange) curve shows the effective PN order of the leading (subleading) pseudo-PN term, i.e. the term with a smaller (greater) $V$ exponent on $\bar{f}$ in Eq. (24), with respect to the polar angle of the latent space $\theta$. The green curve then shows the ratio between the two effective PN coefficients, with the subleading one divided by the leading one. Observe that the subleading PN order never coincides with the leading PN order, i.e. the blue and the orange curves never cross. Observe also that the coefficient ratio is $\gtrsim 1$ (right y-axis) for most of the polar angles between the ppE lines. These two observations suggest that the double-term pseudo-PN expansion learned by our VAE does not reduce to the single-term case. Thus, we verify that we have obtained a npE parameterization that non-trivially extends the ppE framework.
@@ IV.1 Injection-Recovery Setup
All injected signals will consist of GWs generated by quasi-circular, non-spinning BBH systems. In particular, we will consider two prototypical systems: a heavy binary with total mass $M_{\mathrm{inj}}=35\,M_{\odot}$, and a lighter binary with $M_{\mathrm{inj}}=15\,M_{\odot}$, both with a mass ratio of $q_{\mathrm{inj}}=2/3$. These signals will be assumed to have been detected by a Hanford-Livingston-Virgo network, with an O5 design sensitivity [27] [^4] https://dcc.ligo.org/LIGO-T2000012/public. We do not inject the signals into specific realizations of noise because we wish to study the averaged outcome of parameter estimation, which is independent of a given noise artifact. The signals are analyzed in a frequency range between $10\,\mathrm{Hz}$ and $0.018/M_{\mathrm{inj}}$, the latter of which is chosen so that only the inspiral part of the signal is taken into account. For each signal, we scale the luminosity distance such that the network SNR is always fixed at $40$ for the frequency range specified above.
@@ IV.2 Constraining deviations from GR
Let us begin by considering GR injections and an npE model for recovery. Figure 7 shows the npE posteriors in the $(z_{1},z_{2})$ and the $(z_{b},\varphi)$ parameterizations for both the heavy and the light BBHs. For comparison, this figure also shows the ppE posteriors for a fixed set of ppE indices $b_{\mathrm{ppE}}$, which are the same as those used in the npE training set. For each fixed value of $b_{\mathrm{ppE}}$, the $\beta_{\mathrm{ppE}}$ posterior is transformed to the latent $(z_{1},z_{2})$ representation using the npE encoder to allow for direct comparisons.
Figure 7: Predicted constraints on non-GR modifications with the npE framework. GR injections of a heavy BBH (panels (a) and (b)) and a light BBH (panels (c) and (d)) are recovered with an npE model in the $(z_{1},z_{2})$ parameterization (panels (a) and (c)) and the $(z_{b},\varphi)$ parameterization (panels (b) and (d)). The posterior distributions are shown in blue, where the light and dark blue boundaries represent a 90% and 50% credible contours respectively. For reference, ppE models (i.e. with fixed ppE indices) are represented in this figure with gray dotted lines, while regions in the parameter space that are outside of the npE prior are shaded in gray. In order to allow for comparisons between the npE analysis and an independent ppE analysis, we also present the 90% credible region of the $\beta_{\mathrm{ppE}}$ posterior, mapped to the $(z_{1},z_{2})$ or $(z_{b},\varphi)$ latent parameter space with solid orange lines. The insets in each panel shows the npE recovery with a prior range that is smaller than the standard one given by the EFT prior boundary, which is indicated by the limit of the inset axis. Each panel is also accompanied by 1D marginalized posteriors, where the dashed lines mark the 90% credible intervals. Observe that the npE model is able to constrain a wide range of the latent parameter space, with the constraints becoming more stringent in the direction of negative PN order deviations, i.e. the white regions in the panels are excluded at 90% confidence. Higher PN order deviations, however, are constrained less well, as expected. Finally, observe that the npE 90% credible contours are comparable to the ppE 90% credible intervals, when the prior range is properly chosen to resolve the npE contours near the origin. This indicates that the npE model is as good as the ppE one at constraining deviations from GR, with the advantage of being able to explore a wider region of theory parameter space.
@@ IV.3 Detecting Leading-PN-Order Deviations from GR
The resulting boundary locations are summarized in Fig. 9. For comparison, the ppE constraints from Sec. IV.2 are added to the plots. We have checked that the ppE 90% CIs of $\beta_{\mathrm{ppE}}$ critically touch 00 when the modifications are injected at the same level of these constraints. Therefore, we may think of the ppE constraints as an approximation to the ppE detection boundary. Due to the correlation across the npE latent space, the npE detection boundary is always above the ppE one, and the npE resolution boundary is always above the npE detection boundary (if not overlapping). We note that, although the npE test is not as good as the ppE test at detecting a modification, it facilitates resolution of the theory type, which is impossible with a ppE test.
@@ IV.4 Detecting High-PN-Order Deviations from GR
Figure 10: Recovering an EdGB modification with the npE framework. The injected signal is generated with the light BBH, and the EdGB modification is computed to next-to-next-to-leading order with $\sqrt{\alpha_{\mathrm{EdGB,inj}}}=2.5\,\mathrm{km}$. Panels (a) and (b) show the posterior in the non-GR sector with the $z_{1}$-$z_{2}$ parameterization and the $z_{b}$-$\varphi$ parameterization, respectively. The up-pointing (down-pointing) triangle marks the npE representation of the EdGB phase modification calculated to next-to-next-to-leading (leading) order, using the npE encoder. Panel (c) shows the posterior of the mass parameters with the $z_{1}$-$z_{2}$ parameterization. The square dot at the intersection of solid black lines marks the injected values. Other aspects of panels (a)–(c) follow the same format as in Fig. 7. Panel (d) shows the reconstruction error of the total phase (GR plus modification) as a function of frequency. The solid blue (orange) curve is the error of the npE (ppE) reconstruction taking source parameters as given in the injection, and non-GR parameters as given by the up-pointing (down-pointing) triangle in panels (a) and (b). The dashed blue (orange) curve is the error of the npE (ppE) reconstruction taking the median of the npE (ppE) marginalized posterior of each parameter, where the npE posterior is obtained with the $z_{1}$-$z_{2}$ parameterization. The background shading in panel (d) shows the weight of each frequency point in the SNR integral, proportional to $fA(f)^{2}/S_{n}(f)$. Observe that the npE posterior in the non-GR sector finds the theory type as slightly deviated from the leading $-1$PN order, and the npE recovery of the masses is less biased than the ppE one. Although the ppE estimate of the masses is more precise, it is also less accurate because it is completely biased, whereas our npE framework provides a much more accurate result, albeit it less precise.
Let us first take a look at how the npE framework models such a modification. The npE representation of the given EdGB modification, found using the encoder mapping, is shown by the black up-pointing triangles in Fig. 10 (a) and (b). Observe that the npE representation is shifted towards $-0.5$PN order in response to the higher PN-order terms. This is to be compared to the leading-PN order, ppE counterpart, which is shown with down-pointing triangles on the $-1$PN line. Let us then use the npE up-pointing triangle to reconstruct the phase and compare this with the original EdGB injection. The difference, namely the modeling error, is shown by the blue solid curve in Fig. 10 (d). In comparison, the ppE counterpart, obtained with the ppE down-pointing triangle, is shown by the orange solid curve. We note that the npE modeling error is significantly smaller than the ppE one in the entire frequency range, confirming that the npE parameterization captures the higher-PN-order corrections in the EdGB modification much better than the ppE model.
@@ IV.5 Detecting non-PN-like deviations from GR
Although Eq. (34) is developed for binary neutron stars, we apply it to the BBHs that we study in this paper, as a proof of principle. Again, we focus on the light BBH source for better resolution of the inspiral. We assume the dark charge is distributed such that $\alpha_{\mathrm{DP,inj}}=0.05$ and $\gamma_{\mathrm{DP,inj}}=0.02$, which is within the range of validity of Eq. (34[47]. For the dark photon, we consider two injected Compton wavelengths, $\lambda_{\mathrm{DP,inj}}=1400\,\mathrm{km}$ and $300\,\mathrm{km}$. In the first case, the dipole emission is activated at $64\,\mathrm{Hz}$. We have seen in Fig. 10 (d) that this is roughly where most of the SNR is aggregated. In the second case, the dipole emission is activated at $273\,\mathrm{Hz}$ (or $0.020/M_{\mathrm{inj}}$), which is not within the frequency range of the injected signal; in this case, the dark-photon modification is proportional to the $F_{3}$ function in Eq. (34). Moreover, the $\lambda_{\mathrm{DP,inj}}$ value is chosen such that the argument of $F_{3}$ is slightly above unity around $64\,\mathrm{Hz}$. Thus, this scenario features non-perturbative behavior that is encoded by $F_{3}$. We could further reduce $\lambda_{\mathrm{DP,inj}}$ to make $F_{3}$ even more non-perturbative, but doing so would also reduce the overall magnitude of $F_{3}$, which would eventually bring us back to a GR injection.
Figures 11 and 12 show the npE results of the above two injection cases, respectively and allow us to make several observations. First, observe that for both injections, the npE posteriors recover $z_{i}$ very differently from the npE representation of the injected $\Phi_{\mathrm{DP}}$ (the black up-pointing triangles). For the first injection, Fig. 11 (d) indicates that the npE latent space does not cover anything like a dipole activation, as the step function in Eq. (34) is not eliminated in errors reconstructed from either the representation point or the posterior median. For the second injection, Fig. 12 (d) suggests that the npE model covers the non-perturbative $F_{3}$ by the same criteria in Sec. IV.4, as the shape and magnitude of the phase reconstruction errors resemble those in Fig. 10 (d) for EdGB. However, this is incorrect. For the injection, the npE representation point lies right on the $0.5$PN line, so if ppE had been considered and all indices had been explored, the same reconstruction could be reproduced. For the recovery, the npE posterior differs a lot from the representation point, with a majority disfavoring all ppE theories included by the latent space. Moreover, the posterior recovery of the chirp mass fails, as shown in Figs. 11 (c) and 12 (c), leading to a strong bias in chirp mass. All these observations suggest that the npE framework cannot model those non-PN-expandable modified signals more correctly than what the ppE framework can achieve.

### v2

@@ III.6 Training procedure and results
We implement the networks described in Secs. III.4 and III.5 with PyTorch [103]. For training, we choose a combination of the AdamW optimizer [104], with a weight decay of $10^{-4}$, and the ExponentialLR scheduler, with a learning rate decay of by a factor of $0.9$ every epoch [^5] In each epoch, the training set is randomly split into mini-batches. Then minimization of the loss is progressively performed by the optimizer batch by batch so as to evade local minimums. The completion of an epoch marks one full iteration over the entire training set. The learning rate controls how much the network parameters get updated by each minibatch of training data. At the end of each epoch, the learning rate is reduced by the scheduler, so that optimization converges over epochs. See [105] for more details about the related concepts.. We start by training the VAE for $50$ epochs with a batch size of $64$ and an initial learning rate of $10^{-4}$. The coefficient of the KL loss is fixed to $\kappa=10^{-6}$ during the process. Then, we freeze the VAE parameters and train the secondary network for another 50 epochs using the same batch size and initial learning rate. Here, the value of $\kappa$ is empirically determined to give the best learning result for the shape function. The other settings, such as the learning rate and the batch size, are free to vary in some reasonable range, at most impacting the speed of convergence.
@@
Figure 4 shows the training and validation losses as functions of epochs for the VAE [panel (a)] and the secondary network [panel (b)]. Observe that both the VAE and the secondary network reach convergence within the maximum number of epochs explored. The validation loss of the secondary network is considerably higher than the training loss after the loss has converged, implying overfitting. However, as mentioned in Sec. III.5, the secondary network does not necessarily require high accuracy, since it only affects the EFT prior boundary. The validation loss after convergence implies a $\sim 6\%$ fractional error for the modeling of the scale function $T$ [^6] According to Eq. (31), this error is roughly the exponential of the square-root average loss minus one. Here, the average loss is read from the rightmost orange marker in Fig. 4 (b)., which is acceptable for our purposes.
@@
The representation of ppE theories in the latent space (given by the encoder mean) is shown in Fig. 5 (a). These ppE theories map to radial lines, which is not surprising given the npE prescription in Eqs. (14) and (15), and the fact that ppE phases associated with the same $b_{\mathrm{ppE}}$, do not vary in shape. What matters is that these ppE theories are well organized in the latent space. In particular, the mapping is injective, in the sense that there is a one-to-one relation between the ppE index $b_{\mathrm{ppE}}$ (or PN order) and the polar angle (up to $\pi$) of the latent line, so each theory maps to one line and different theories separate. Moreover, the mapping is ordered, in the sense that theories with deformations that enter at similar PN orders are also close to each other in angle in the latent space. We highlight that we arrive at this result in a completely unsupervised manner, i.e., ordering is not imposed explicitly.
@@
Figure 5 (a) also presents ppE theories that introduce modifications that enter at half-integer PN orders. These theories are not included in the modified gravity dataset for training, but they are mapped to the latent space in a similar way to those in the training set. This means that the VAE understands well the structure of ppE theories in the latent space. These observations are further confirmed by Fig. 5 (b), where we present the shape functions generated from a sample of latent space angles, which is much denser than the distribution of angles of the ppE theories contained in the training set. These functions form a band when overlaid, suggesting that the ppE parametrization has been continuously extended by the npE latent space.
As discussed in Sec. III.4, the pseudo-PN expansion adopted by the decoder leads to a trivial extension of the ppE parametrization when $N_{p}=1$. Although we have chosen $N_{p}=2$, the two terms in Eq. (24) may lead to degeneracies after training, and reduce to the trivial case. Therefore, we must investigate the behavior of the learned pseudo-PN expansion across the latent space, which is illustrated in Fig. 6. In this figure, the blue (orange) curve shows the effective PN order of the leading (subleading) pseudo-PN term, i.e. the term with a smaller (greater) $V$ exponent on $\bar{f}$ in Eq. (24), with respect to the polar angle of the latent space $\theta$. The green curve then shows the ratio between the two effective PN coefficients, with the subleading one divided by the leading one. Observe that the subleading PN order never coincides with the leading PN order, i.e. the blue and the orange curves never cross. Observe also that the coefficient ratio is $\gtrsim 1$ (right y-axis) for most of the polar angles between the ppE lines. These two observations suggest that the double-term pseudo-PN expansion learned by our VAE does not reduce to the single-term case. Thus, we verify that we have obtained a npE parametrization that nontrivially extends the ppE framework.
@@ IV.1 Injection-recovery setup
All injected signals will consist of GWs generated by quasicircular, nonspinning BBH systems. In particular, we will consider two prototypical systems: a heavy binary with total mass $M_{\mathrm{inj}}=35\,M_{\odot}$, and a lighter binary with $M_{\mathrm{inj}}=15\,M_{\odot}$, both with a mass ratio of $q_{\mathrm{inj}}=2/3$. These signals will be assumed to have been detected by a Hanford-Livingston-Virgo network, with an O5 design sensitivity [27] [^7] https://dcc.ligo.org/LIGO-T2000012/public. We do not inject the signals into specific realizations of noise because we wish to study the averaged outcome of parameter estimation, which is independent of a given noise artifact. The signals are analyzed in a frequency range between $10\,\mathrm{Hz}$ and $0.018/M_{\mathrm{inj}}$, the latter of which is chosen so that only the inspiral part of the signal is taken into account. For each signal, we scale the luminosity distance such that the network SNR is always fixed at $40$ for the frequency range specified above.
@@ IV.2 Constraining deviations from GR
Let us begin by considering GR injections and an npE model for recovery. Figure 7 shows the npE posteriors in the $(z_{1},z_{2})$ and the $(z_{b},\varphi)$ parametrizations for both the heavy and the light BBHs. For comparison, this figure also shows the ppE posteriors for a fixed set of ppE indices $b_{\mathrm{ppE}}$, which are the same as those used in the npE training set. For each fixed value of $b_{\mathrm{ppE}}$, the $\beta_{\mathrm{ppE}}$ posterior is transformed to the latent $(z_{1},z_{2})$ representation using the npE encoder to allow for direct comparisons.
Figure 7: Predicted constraints on non-GR modifications with the npE framework. GR injections of a heavy BBH [panels (a) and (b)] and a light BBH [panels (c) and (d)] are recovered with an npE model in the $(z_{1},z_{2})$ parametrization [panels (a) and (c)] and the $(z_{b},\varphi)$ parametrization [panels (b) and (d)]. The posterior distributions are shown in blue, where the light and dark blue boundaries represent a 90% and 50% credible contours respectively. For reference, ppE models (i.e., with fixed ppE indices) are represented in this figure with gray dotted lines, while regions in the parameter space that are outside of the npE prior are shaded in gray. In order to allow for comparisons between the npE analysis and an independent ppE analysis, we also present the 90% credible region of the $\beta_{\mathrm{ppE}}$ posterior, mapped to the $(z_{1},z_{2})$ or $(z_{b},\varphi)$ latent parameter space with solid orange lines. The insets in each panel shows the npE recovery with a prior range that is smaller than the standard one given by the EFT prior boundary, which is indicated by the limit of the inset axis. Each panel is also accompanied by 1D marginalized posteriors, where the dashed lines mark the 90% credible intervals. Observe that the npE model is able to constrain a wide range of the latent parameter space, with the constraints becoming more stringent in the direction of negative PN order deviations, i.e. the white regions in the panels are excluded at 90% confidence. Higher PN order deviations, however, are constrained less well, as expected. Finally, observe that the npE 90% credible contours are comparable to the ppE 90% credible intervals, when the prior range is properly chosen to resolve the npE contours near the origin. This indicates that the npE model is as good as the ppE one at constraining deviations from GR, with the advantage of being able to explore a wider region of theory parameter space.
@@ IV.3 Detecting leading-PN-order deviations from GR
The resulting boundary locations are summarized in Fig. 9. For comparison, the ppE constraints from Sec. IV.2 are added to the plots. We have checked that the ppE 90% CIs of $\beta_{\mathrm{ppE}}$ critically touch 00 when the modifications are injected at the same level of these constraints. Therefore, we may think of the ppE constraints as an approximation to the ppE detection boundary. Due to the correlation across the npE latent space, the npE detection boundary is always above the ppE one, and the npE resolution boundary is always above the npE detection boundary (if not overlapping). We note that, although the npE test is not as good as the ppE test at detecting a modification, it facilitates resolution of the theory type, which is impossible with a ppE test. Moreover, the boundaries in Fig. 9 are not constant limits as they should scale by $1/\mathrm{SNR}$. With that in mind, future observation of high-SNR events will make it possible to detect smaller non-GR deviations.
@@ IV.4 Detecting high-PN-order deviations from GR
Figure 10: Recovering an EdGB modification with the npE framework. The injected signal is generated with the light BBH, and the EdGB modification is computed to next-to-next-to-leading order with $\sqrt{\alpha_{\mathrm{EdGB,inj}}}=2.5\,\mathrm{km}$. Panels (a) and (b) show the posterior in the non-GR sector with the $z_{1}$-$z_{2}$ parametrization and the $z_{b}$-$\varphi$ parametrization, respectively. The up-pointing (down-pointing) triangle marks the npE representation of the EdGB phase modification calculated to next-to-next-to-leading (leading) order, using the npE encoder. Panel (c) shows the posterior of the mass parameters with the $z_{1}$-$z_{2}$ parametrization. The square dot at the intersection of solid black lines marks the injected values. Other aspects of panels (a)–(c) follow the same format as in Fig. 7. Panel (d) shows the reconstruction error of the total phase (GR plus modification) as a function of frequency. The solid blue (orange) curve is the error of the npE (ppE) reconstruction taking source parameters as given in the injection, and non-GR parameters as given by the up-pointing (down-pointing) triangle in panels (a) and (b). The dashed blue (orange) curve is the error of the npE (ppE) reconstruction taking the median of the npE (ppE) marginalized posterior of each parameter, where the npE posterior is obtained with the $z_{1}$-$z_{2}$ parametrization. The background shading in panel (d) shows the weight of each frequency point in the SNR integral, proportional to $fA(f)^{2}/S_{n}(f)$. Observe that the npE posterior in the non-GR sector finds the theory type as slightly deviated from the leading $-1$PN order, and the npE recovery of the masses is less biased than the ppE one. Although the ppE estimate of the masses is more precise, it is also less accurate because it is completely biased, whereas our npE framework provides a much more accurate result, albeit it less precise.
Let us first take a look at how the npE framework models such a modification. The npE representation of the given EdGB modification, found using the encoder mapping, is shown by the black up-pointing triangles in Fig. 10 (a) and (b). Observe that the npE representation is shifted toward $-0.5$PN order in response to the higher PN-order terms. This is to be compared to the leading-PN order, ppE counterpart, which is shown with down-pointing triangles on the $-1$PN line. Let us then use the npE up-pointing triangle to reconstruct the phase and compare this with the original EdGB injection. The difference, namely the modeling error, is shown by the blue solid curve in Fig. 10 (d). In comparison, the ppE counterpart, obtained with the ppE down-pointing triangle, is shown by the orange solid curve. We note that the npE modeling error is significantly smaller than the ppE one in the entire frequency range, confirming that the npE parametrization captures the higher-PN-order corrections in the EdGB modification much better than the ppE model.
@@ IV.5 Detecting non-PN-like deviations from GR
Although Eq. (34) is developed for binary neutron stars, we apply it to the BBHs that we study in this paper, as a proof of principle. Again, we focus on the light BBH source for better resolution of the inspiral. We assume the dark charge is distributed such that $\alpha_{\mathrm{DP,inj}}=0.05$ and $\gamma_{\mathrm{DP,inj}}=0.02$, which is within the range of validity of Eq. (34[49]. For the dark photon, we consider two injected Compton wavelengths, $\lambda_{\mathrm{DP,inj}}=1400\,\mathrm{km}$ and $300\,\mathrm{km}$. In the first case, the dipole emission is activated at $64\,\mathrm{Hz}$. We have seen in Fig. 10 (d) that this is roughly where most of the SNR is aggregated. In the second case, the dipole emission is activated at $273\,\mathrm{Hz}$ (or $0.020/M_{\mathrm{inj}}$), which is not within the frequency range of the injected signal; in this case, the dark-photon modification is proportional to the $F_{3}$ function in Eq. (34). Moreover, the $\lambda_{\mathrm{DP,inj}}$ value is chosen such that the argument of $F_{3}$ is slightly above unity around $64\,\mathrm{Hz}$. Thus, this scenario features nonperturbative behavior that is encoded by $F_{3}$. We could further reduce $\lambda_{\mathrm{DP,inj}}$ to make $F_{3}$ even more nonperturbative, but doing so would also reduce the overall magnitude of $F_{3}$, which would eventually bring us back to a GR injection.
Figures 11 and 12 show the npE results of the above two injection cases, respectively and allow us to make several observations. First, observe that for both injections, the npE posteriors recover $z_{i}$ very differently from the npE representation of the injected $\Phi_{\mathrm{DP}}$ (the black up-pointing triangles). For the first injection, Fig. 11 (d) indicates that the npE latent space does not cover anything like a dipole activation, as the step function in Eq. (34) is not eliminated in errors reconstructed from either the representation point or the posterior median. For the second injection, Fig. 12 (d) suggests that the npE model covers the nonperturbative $F_{3}$ by the same criteria in Sec. IV.4, as the shape and magnitude of the phase reconstruction errors resemble those in Fig. 10 (d) for EdGB. However, this is incorrect. For the injection, the npE representation point lies right on the $0.5$PN line, so if ppE had been considered and all indices had been explored, the same reconstruction could be reproduced. For the recovery, the npE posterior differs a lot from the representation point, with a majority disfavoring all ppE theories included by the latent space. Moreover, the posterior recovery of the chirp mass fails, as shown in Figs. 11 (c) and 12 (c), leading to a strong bias in chirp mass. All these observations suggest that the npE framework cannot model those non-PN-expandable modified signals more correctly than what the ppE framework can achieve.